Consider this initial-rate data at a certain temperature for the reaction described by

2NO2(g)+O3(g) ----> N2O5(g) + O2(g)

[NO2]0 (M)         [O3]0(M)      Initial rate (M/s)

0.650                  0.800           2.47x10^4

1.10                    0.800           4.18x10^4

1.76                    1.40             11.70x10^4

Determine the value and units of the rate constant.

K=_____
Answer #1
Concepts and reason

This problem is based on the concept of chemical kinetics.

The rate constant of a reaction is determined by rate of the reaction. For this, first order of the reaction is calculated.

Fundamentals

For the reaction,

A+BC+D{\rm{A}} + {\rm{B}} \to {\rm{C}} + {\rm{D}}

The rate of the reaction is written as follows:

r=k[A]x[B]yr = k{\left[ A \right]^x}{\left[ B \right]^y}

Here, [A]\left[ A \right] , [B]\left[ B \right] are the molar concentration of reactants A and B respectively. And x and y are order of the reaction with respect to A and B.

The reaction is as follows:

2NO2(g)+O3(g)N2O5(g)+O2(g){\rm{2N}}{{\rm{O}}_2}\left( g \right) + {{\rm{O}}_3}\left( g \right) \to {{\rm{N}}_2}{{\rm{O}}_5}\left( g \right) + {{\rm{O}}_2}\left( g \right)

The rate of the reaction is written as follows:

r=k[NO2]x[O3]yr = k{\left[ {{\rm{N}}{{\rm{O}}_2}} \right]^x}{\left[ {{{\rm{O}}_3}} \right]^y}

The order of reaction is calculated by dividing the two rate law equations as follows:

2.47×1044.18×104=k[NO2]x[O3]yk[NO2]x[O3]y\frac{{2.47 \times {{10}^4}}}{{4.18 \times {{10}^4}}} = \frac{{k{{\left[ {{\rm{N}}{{\rm{O}}_2}} \right]}^x}{{\left[ {{{\rm{O}}_3}} \right]}^y}}}{{k{{\left[ {{\rm{N}}{{\rm{O}}_2}} \right]}^x}{{\left[ {{{\rm{O}}_3}} \right]}^y}}}

Here, x is order with respect to concentration of NO2{\rm{N}}{{\rm{O}}_2} , y is order with respect to concentration of O3{{\rm{O}}_3} .

Now, according to the table, the order of reaction is calculated as follows:

2.47×1044.18×104=k[0.65]x[0.8]yk[1.1]x[0.8]y0.59=[0.59]xx=1\begin{array}{c}\\\frac{{2.47 \times {{10}^4}}}{{4.18 \times {{10}^4}}} = \frac{{k{{\left[ {0.65} \right]}^x}{{\left[ {0.8} \right]}^y}}}{{k{{\left[ {1.1} \right]}^x}{{\left[ {0.8} \right]}^y}}}\\\\0.59 = {\left[ {0.59} \right]^x}\\\\x = 1\\\end{array}

Thus, the order of reaction with respect to NO2{\rm{N}}{{\rm{O}}_2} is 1.

Also, Now divide the equations as follows:

2.47×10411.7×104=k[0.65]1[0.8]yk[1.76]1[1.4]y[0.571]1=[0.571]yy=1\begin{array}{c}\\\frac{{2.47 \times {{10}^4}}}{{11.7 \times {{10}^4}}} = \frac{{k{{\left[ {0.65} \right]}^1}{{\left[ {0.8} \right]}^y}}}{{k{{\left[ {1.76} \right]}^1}{{\left[ {1.4} \right]}^y}}}\\\\{\left[ {0.571} \right]^1} = {\left[ {0.571} \right]^y}\\\\y = 1\\\end{array}

Thus, the order of reaction with respect to O3{{\rm{O}}_3} is 1.

Thus, rate law equation is written as follows:

r=k[NO2]1[O3]1r = k{\left[ {{\rm{N}}{{\rm{O}}_2}} \right]^1}{\left[ {{{\rm{O}}_3}} \right]^1}

Rearrange for the rate constant as follows:

k=r[NO2]1[O3]1k = \frac{r}{{{{\left[ {{\rm{N}}{{\rm{O}}_2}} \right]}^1}{{\left[ {{{\rm{O}}_3}} \right]}^1}}}

Now, substitute 11.7×10411.7 \times {10^4} for r , 1.76 for [NO2]\left[ {{\rm{N}}{{\rm{O}}_2}} \right] and 1.40 for [O3]\left[ {{{\rm{O}}_3}} \right] .

k=11.7×104(1.76)(1.40)=4.74×104mol1Lsec1\begin{array}{c}\\k = \frac{{11.7 \times {{10}^4}}}{{\left( {1.76} \right)\left( {1.40} \right)}}\\\\ = 4.74 \times {10^4}{\rm{ mo}}{{\rm{l}}^{ - 1}}{\rm{ L se}}{{\rm{c}}^{ - 1}}\\\end{array}

Ans:

The rate constant is 4.74×104mol1Lsec14.74 \times {10^4}{\rm{ mo}}{{\rm{l}}^{ - 1}}{\rm{ L se}}{{\rm{c}}^{ - 1}} .